Project

General

Profile

NDOS CommissionMtng 20101202

Attendance: Jon, Leon, Ryan, Athans, Nathan, Peter, Andrew, Kurt, Mark Messier, Mark Mathis, Minerba, Greg, Art, Alena, Enrique, Tricia, Gavin, Denis, Mat, Carl B., Robert H.

Agenda:
  1. DAQ/Electronics Update & Plans
  2. DCS Status & Plans
  3. Shifters' report
  4. AOB

DAQ/Electronics Update & Plans:

  • Denis and Mat have been collecting data in DSO, and recently began collecting data with different APD thresholds.
  • Recently, some pixel masks have been determined from the data collected in DSO mode and applied to "normal" runs (10 Hz calibration trigger). The results have looked very promising with some very nice runs taken overnight.
  • Peter is investigating an anomalous 80 us trigger timing offset with the NuMI spill trigger.
  • Work has been going on over at FCC on throttling of data in DCMs
  • APD Cooling:
    • Chiller is broken, we are waiting on a replacement part (valve). That part should be on its way.
    • Leon: another issue that needs to be resolved before we can consider cooling the APDs is that the integration of TEC control by the FEB in case of certain failure modes. Nathan said this is #3 or #4 in his queue.
    • The plan is to continue collecting data in both DSO mode as well as regular runs overnight with different APD thresholds.

DCS Status & Plans

  • Athans is getting caught up on the DCS requirements.
  • Athans will be here next week to work on a Labview program to read out the DCS data from the FEBs.

Shifters' Report

  • Robert H. reported that he had been RTV'ing block 6.
  • Both he and Carl B. have been taking some notes regarding shifter/DAQ documentation and will get those the Jon ASAP.

AOB

  • Andrew showed some analysis of some DSO data collected recently. See doc-db 5409 for details, but some important points:
    • Pedestals in general look pretty good, with typical widths of 15-30 ADC counts at 400V bias and 5-8 ADC counts at min. bias voltages (~5V)
    • Pedestal noise appears to be unstructured.
    • Nearly all problem (noisy) channels appear to be pixels at the corners (or along the edges) of the APDs. This noise appears to occasionally induce noise on neighboring pixels.
    • We can mask out the noisy channels; this eliminates ~50% of all channels in the IPND.